# When does bottom-up beat top-down in hierarchical community detection?

### v2

@@ A.2.3 Proof of Proposition 1
When $h_{ac}>0$, by substituting $\zeta(h)$ with Equation (3.2), the left hand side of Condition (3) can be written as
@@
When $h_{ac}=0$, we can replace $\zeta(h)$ with Equation (3.2) and denote $\frac{\eta}{2^{d-1}}$ by $\zeta$ for convenience. This allows us to express the left-hand side of Condition (3) as follows:
@@
Notice that $\eta_{+}=\frac{\sqrt{\bar{p}_{1}-p_{d-1}}(2\sqrt{\bar{p}_{1}-p_{0}}-\sqrt{\bar{p}_{1}-p_{d-1}})}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}+\frac{1}{2}>\frac{1}{2}$, and hence the condition $\eta>\eta_{+}$ cannot be verified (recall that $20\leq\eta<1/$ 2). In contrast, we have $\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}-\frac{1}{2}=\frac{p_{d-1}-\bar{p}_{1}}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}\leq 0$ and hence $\eta_{-}\leq\frac{1}{2}$. Moreover, using
@@ A.2.4 Proof of Lemma 3
$\mathbb{E}w\left(O_{ka},O_{\ell c}\right)\ =\ \left(\frac{N}{K}\right)^{2}\times\begin{cases}\zeta(h_{1})\zeta(h_{2})p_{h_{1}}&\text{ if }0\leq h_{1}<h_{ac}\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(h_{ac})\zeta(h_{2})\bar{p}_{h_{ac}+1}&\text{ if }h_{1}=h_{ac}\text{ and }h_{ac}<h_{2}\leq d,\\ \zeta(h_{1})\zeta(h_{2})p_{h_{ac}}&\text{ if }h_{ac}<h_{1}<h_{2}\leq d,\\ \zeta^{2}(h_{1})\bar{p}_{h_{1}+1}&\text{ if }0\leq h_{1}=h_{2}<h_{ac},\\ \zeta^{2}(h_{1})p_{h_{ac}}&\text{ if }h_{ac}\leq h_{1}=h_{2}\leq d,\\ \end{cases}$
@@
$\mathbb{E}w\left(O_{ka},O_{\ell b}\right)\ =\ \left(\frac{N}{K}\right)^{2}\times\begin{cases}\zeta(h_{1})\zeta(h_{2})p_{h_{1}}&\text{ if }0\leq h_{1}<d-1\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(d-1)\zeta(d)p_{d}&\text{ if }(h_{1},h_{2})=(d-1,d),\\ \zeta^{2}(h_{1})\bar{p}_{h_{1}+1}&\text{ if }h_{1}=h_{2}<d-1,\\ \zeta^{2}(h_{1})p_{d-1}&\text{ if }d-1\leq h_{1}=h_{2}\leq d,\\ \end{cases}$
@@
$\left(\frac{N}{K}\right)^{2}\times\begin{cases}0&\text{ if }0\leq h_{1}<h_{ac}\text{ and }h_{1}<h_{2}\leq d,\\ \zeta(h_{ac})\zeta(h_{2})\left(p_{h_{ac}}-\bar{p}_{h_{ac}+1}\right)&\text{ if }h_{1}=h_{ac}<h_{2}\leq d,\\ \zeta(h_{1})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)&\text{ if }h_{ac}<h_{1}<h_{2}\leq d\text{ and }h_{1}<d-1,\\ \zeta(d-1)\zeta(d)\left(p_{d}-p_{h_{ac}}\right)&\text{ if }(h_{1},h_{2})=(d-1,d),\\ 0&\text{ if }0\leq h_{1}=h_{2}<h_{ac},\\ \zeta^{2}(h_{1})\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)&\text{ if }h_{ac}\leq h_{1}=h_{2}<d-1,\\ \zeta^{2}(h_{1})\left(p_{d-1}-p_{h_{ac}}\right)&\text{ if }d-1\leq h_{1}=h_{2}\leq d.\\ \end{cases}$
@@ A.2.5 Additional computation for the proof of Lemma 3
$\Delta\rho\ =$ $(1+o(1))\Bigg{(}-\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{ac})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(\bar{p}_{h_{ac}+1}-p_{h_{ac}}\right)$
@@
[…]
@@
$\Delta\rho\ =$ $(1+o(1))\Bigg{(}-\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
@@
[…]
@@ B.1 Proof of Theorem 3
In the following, we denote by $\mathcal{A}=\{A_{1},\cdots,A_{t}\}$ a collection of $t$ non-empty and non-overlapping subsets of $[K]$ such that $\cup_{r=1}^{t}A_{r}=[K]$. An algorithm exactly recovers $\mathcal{A}$ if it assigns each node $i$ in $G$ to an element of $\{A_{1},\cdots,A_{t}\}$ that contains its true community (up to a relabelling of the $A_{r}$’s) with probability $1-o(1)$. For two non-overlapping subsets $A,B\subset[K]$, we denote by $\mathrm{CH}\left(A,B\right)=\mathrm{CH}(A,B,\pi,p)$ the quantity
@@
    the agnostic-degree-profiling algorithm of [3] exactly recovers the finest partition with threshold $\tau=(1+\Omega(1))\,\,N^{-1}\log N$.
@@
The proof of point (i) can be found in [2, Theorem 1] while point (ii) corresponds to [3, Theorem 4]. ∎
@@
Conversely, point (ii) of Theorem 4 implies that we can recover the finest partition $q^{*}$. In the case of an HSBM, this corresponds to the largest $q$ such that $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. Since $q\mapsto I_{q}$ is non-decreasing we have $q\geq q^{*}$. Moreover, using Lemma 4, we proceed as in the proof of Theorem 1 to show that we recover the tree $\mathcal{T}[\mathcal{S}_{\leq q^{*}}]$. In particular, we can recover the super-communities at any higher depth $q\geq q^{*}$, which are exactly the depths verifying $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. ∎
@@ B.2 Proof of Lemma 1
$\left|\Gamma_{a,b}(v,w)\right|\ =\ \begin{cases}1&\text{ if }\mathrm{sim}(a,v)=d\text{ and }\mathrm{sim}(b,w)=s\text{ (this is equivalent to $v=a$ and $w=u$)},\\ 1&\text{ if }\mathrm{sim}(a,v)=s\text{ and }\mathrm{sim}(b,w)=d\text{ (this isequivalent to $v=u$ and $w=b$)},\\ 2^{k-1}&\text{ if }\mathrm{sim}(a,v)=d-k\text{ and }\mathrm{sim}(b,w)=s\text{ for some }k\in[d-s-1],\\ 2^{k-1}&\text{ if }\mathrm{sim}(a,v)=s\text{ and }\mathrm{sim}(b,w)=d-k\text{ for some }k\in[d-s-1],\\ 0&\text{ otherwise.}\end{cases}$
@@
Finally, let $P,Q$ be two probability distributions. By the concavity of $t\mapsto(1-t)\operatorname{D}_{t}(P\|Q)$ and the relation $(1-t)\operatorname{D}_{t}(P\|Q)=t\operatorname{D}_{1-t}(Q\|P)$ (see for example [34]), the function inside the $\sup$ of Equation B.1 is concave and symmetric around $t=1/2$. Therefore the $\sup$ of Equation B.1 is achieved at $t=1/2$. Hence,
@@ B.3 Comparing top-down versus bottom-up conditions
$D^{2}-E^{2}\ =$ $\left(\sum_{k=1}^{d-q}2^{k-1}a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}-2\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d}}\sqrt{a_{d-k}}\right)$
[…]
@@ C.1 Synthetic data sets
C.1.1 Unbalanced HSBM
We also examine the accuracy of HCD algorithms on HSBM whose binary tree is not necessarily full and balanced. Similarly to the BTSBM, we still suppose that the depth of the tree $\mathcal{T}$ determines the link probabilities, i.e., $p(u)=p_{|u|}$ for all $u\in\mathcal{T}$. Moreover, we assume that the primitive communities have the same number of nodes. Figures 9 amd 10 compare the different HCD algorithms on HSBM whose respective unbalanced trees are given in Figures 9(a) and 10(a). We observe that bottom-up and top-down perform similarly well, but one sometimes outperforms the other depending on $\beta$ and the metric used. The relative error for the estimation of the tree similarity matrix is defined as $\|S(\widehat{\mathcal{T}},\widehat{\mathcal{C}})-S(\mathcal{T},\mathcal{C})\|^{2}_{F}/\|S(\mathcal{T},\mathcal{C})\|^{2}_{F}$, where the tree-similarity matrix $S\left(\mathcal{T},\mathcal{C}\right)$ is defined in (6.1).
@@ C.1.1 Unbalanced HSBM
Figure 9: (a) Hierarchical tree. (b)-(d) Results of the different HCD algorithms on unbalanced HSBMs whose hierarchical tree is given in Figure 9(a), and with $N=3200$ nodes and $p_{k}=64\cdot\beta^{5-k}\cdot\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars (vertical solid lines) show the standard error but are typically smaller than the symbols.
@@
Figure 10: (a) Hierarchical tree. (b)-(d) Results of the different HCD algorithms on unbalanced BTSBMs whose tree shape is given in Figure 10(a) and with $N=3200$ nodes and $p_{k}=144\cdot\beta^{5-k}\cdot\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars (vertical solid lines) show the standard error but are typically smaller than the symbols.
@@ C.2.1 Military inter-alliance
We next consider the network of military alliances between countries. The data is provided by the Alliance Treaty Obligations and Provisions (ATOP) project [22]. We select the year 2018 (as this is the most recent year available). We define two countries as allied if they share a defensive alliance [^12] In particular, we do not consider non-aggression pacts, as those are more numerous and historically not necessarily well respected.. This leads to a network of 133 countries and 1391 alliances. Some important countries such as India or Switzerland are missing as they do not share any defensive alliances with anybody. Moreover, the graph is not connected, as a small component made of three countries (China, North Korea and Cuba) is disconnected from the rest of the world.
Figures 11 and 12 show the output of bottom-up and top-down algorithms. Bottom-up predicts 7 bottom communities, which represent geopolitical alliances based on political affiliation and geography (European countries, Eurasian countries, Arabic countries, Western African countries and Central/Southern African countries). The top level splits the graph’s largest connected component into 3 clusters: Western countries, Eurasian countries, and African and Middle-East countries. While some of these clusters are also recovered by top-down HCD, the separation of African countries by top-down algorithm appears worse.
@@ C.2.2 Football data set
We also test HCD algorithms on the United States college (American) football dataset [16]. This network represents the schedule of Division I games for the 2000 season. Each node in the network corresponds to a college football team, and edges represent the regular-season games between the teams. The teams are categorized into 11 conferences, in which games are more frequent between the members. Each conference has 8 to 12 teams. We exclude the ”independent” teams which do not belong to any conferences. Since the original community labels appear to be based on the 2001 season, while the edges represent the games played during the 2000 season, we proceed to the same correction as in [13].

### v3

@@ A.2.3 An Adversarial Setting where $\zeta$ is not Non-decreasing
When $h_{ac}>0$, by substituting $\zeta(h)$ with Equation (A.8), the left hand side of Condition (3) can be written as
@@
When $h_{ac}=0$, we can replace $\zeta(h)$ with Equation (A.8) and denote $\frac{\eta}{2^{d-1}}$ by $\zeta$ for convenience. This allows us to express the left-hand side of Condition (3) as follows:
@@
Notice that $\eta_{+}=\frac{\sqrt{\bar{p}_{1}-p_{d-1}}(2\sqrt{\bar{p}_{1}-p_{0}}-\sqrt{\bar{p}_{1}-p_{d-1}})}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}+\frac{1}{2}>\frac{1}{2}$, and hence the condition $\eta>\eta_{+}$ cannot be verified (recall that $0\leq\eta<1/2$). In contrast, we have $\frac{p_{d-1}+\bar{p}_{1}-2p_{0}}{p_{d-1}+3\bar{p}_{1}-4p_{0}}-\frac{1}{2}=\frac{p_{d-1}-\bar{p}_{1}}{2(p_{d-1}+3\bar{p}_{1}-4p_{0})}\leq 0$ and hence $\eta_{-}\leq\frac{1}{2}$. Moreover, using
@@ A.2.4 Proof of Lemma 3
[…]
@@
[…]
@@
[…]
@@ A.2.5 Additional Computation for the Proof of Lemma 3
$\Delta\rho\ =\$$(1+o(1))\Bigg(-\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{ac})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(\bar{p}_{h_{ac}+1}-p_{h_{ac}}\right)$
@@
$\ =\$$(1+o(1))\Bigg(-\sum_{h_{1}=h_{ac}+1}^{d}\boldsymbol{\sum_{h_{2}=h_{ac}+1}^{d}}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(\boldsymbol{p_{h_{1}}}-p_{h_{ac}}\right)$
@@
$\Delta\rho\ =\$$(1+o(1))\Bigg(-\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)$
@@
$\ =\$$(1+o(1))\Bigg(-\underbrace{\sum_{h_{1}=h_{ac}+1}^{d}\sum_{h_{2}=h_{ac}+1}^{d}2B(h_{1})B(h_{2})\zeta(h_{ac})\zeta(h_{2})\left(p_{h_{1}}-p_{h_{ac}}\right)}_{T_{1}}$
@@ B.1 Proof of Theorem 3
In the following, we denote by $\mathcal{A}=\{A_{1},\cdots,A_{t}\}$ a collection of $t$ non-empty and non-overlapping subsets of $[K]$ such that $\cup_{r=1}^{t}A_{r}=[K]$. An algorithm exactly recovers $\mathcal{A}$ if it assigns each node $i$ in $G$ to an element of $\{A_{1},\cdots,A_{t}\}$ that contains its true community (up to a relabelling of the $A_{r}$’s) with probability 1o(1)1-o(1). For two non-overlapping subsets $A,B\subset[K]$, we denote by $\mathrm{CH}\left(A,B\right)=\mathrm{CH}(A,B,\pi,p)$ the quantity
@@
    the agnostic-degree-profiling algorithm of [abbe2015recovering] exactly recovers the finest partition with threshold $\tau=(1+\Omega(1))\,\,N^{-1}\log N$.
@@
The proof of point (i) can be found in [abbe2015community, Theorem 1] while point (ii) corresponds to [abbe2015recovering, Theorem 4]. ∎
@@
Conversely, point (ii) of Theorem 4 implies that we can recover the finest partition $q^{*}$. In the case of an HSBM, this corresponds to the largest $q$ such that $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. Since $q\mapsto I_{q}$ is non-decreasing we have $q\geq q^{*}$. Moreover, using Lemma 4, we proceed as in the proof of Theorem 1 to show that we recover the tree $\mathcal{T}[\mathcal{S}_{\leq q^{*}}]$. In particular, we can recover all the super-communities at any higher depth $q\geq q^{*}$, which are exactly the depths verifying $I_{q}>(1+\Omega(1))\,N^{-1}\log N$. ∎
@@ B.2 Proof of Lemma 1
[…]
@@
Finally, let $P,Q$ be two probability distributions. By the concavity of $t\mapsto(1-t)\operatorname{D}_{t}(P\|Q)$ and the relation $(1-t)\operatorname{D}_{t}(P\|Q)=t\operatorname{D}_{1-t}(Q\|P)$ (see for example [vanErven2014renyi]), the function inside the $\sup$ of Equation B.1 is concave and symmetric around $t=1/2$. Therefore the $\sup$ of Equation B.1 is achieved at $t=1/2$. Hence,
@@ B.3 Comparing Top-down versus Bottom-up Conditions
$D^{2}-E^{2}\ =\$$\left(\sum_{k=1}^{d-q}2^{k-1}a_{d}+\sum_{k=1}^{d-q}2^{k-1}a_{d-k}-2\sum_{k=1}^{d-q}2^{k-1}\sqrt{a_{d}}\sqrt{a_{d-k}}\right)$
$\qquad+\left(\sum_{k=1}^{d-q}\sum_{\begin{subarray}{c}l=1\\ l\neq k\end{subarray}}^{d-q}2^{k+l-2}a_{d-k}-2\sum_{k=1}^{d-q-1}\sum_{l>k}^{d-q}2^{k+l-2}\sqrt{a_{d-k}}\sqrt{a_{d-l}}\right)$
@@ C.1 Synthetic Data Sets
C.1.1 Ternary Tree SBMs
We first show in Figure 7 the dendrograms and trees obtained by top-down and bottom-up algorithms. Although both algorithms generate binary trees, ternary structures appear in Figures 7(b) and 7(c). Nevertheless, we observe in Figure 7(c) that the dendrograms obtained by the top-down algorithm show some inversions.
(a) Ground truth tree
(b) bottom-up
(c) top-down
C.1.2 Unbalanced HSBM
We evaluate the performance of HCD algorithms on HSBM whose binary tree is not necessarily full and balanced. Similar to the BTSBM, we assume that the depth of the tree $\mathcal{T}$ determines the link probabilities, i.e., $p(u)=p_{|u|}$ for all $u\in\mathcal{T}$. Because the tree is unbalanced, the bottom communities no longer have the same depth. In our experiments, the size of a bottom community having depth $k$ is $100\cdot 2^{5-k}$, so that the total number of nodes is $N=3200$.
To assess the accuracy of tree recovery, we define the similarity matrix $S(\mathcal{T},\mathcal{C})$ the NN-by-NN matrix such that for $i\in C_{a}$ and $j\in C_{b}$ (with $a,b\in\mathcal{L}_{\mathcal{T}}$) we have
$\Big(S\left(\mathcal{T},\mathcal{C}\right)\Big)_{ij}\ =\ |\mathrm{lca}_{\mathcal{T}}(a,b)|,$
where $|\mathrm{lca}_{\mathcal{T}}(a,b)|$ is the depth of the lowest common ancestor of $C_{a}$ and $C_{b}$ in the tree $\mathcal{T}$. The tree recovery error is then defined as
$\frac{\|S(\widehat{\mathcal{T}},\widehat{\mathcal{C}})-S(\mathcal{T},\mathcal{C})\|^{2}_{F}}{\|S(\mathcal{T},\mathcal{C})\|^{2}_{F}}.$
This metric quantifies the discrepancy between the estimated and true similarity matrices.
Figures 9 and 10 compare different HCD algorithms on HSBMs whose corresponding unbalanced trees are shown in Figures 9(a) and 10(a). We observe that bottom-up and top-down approaches perform similarly well, though one may outperform the other depending on $\beta$ and the chosen evaluation metric.
@@ C.1.2 Unbalanced HSBM
Figure 9: Performance of top-down and bottom-up algorithms on HSBMs, where the tree is given in Figure 9(a), with $N=3200$ nodes and $p_{k}=64\,\beta^{5-k}\,\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars show the standard error but are typically smaller than the symbols.
@@
Figure 10: Performance of top-down and bottom-up algorithms on HSBMs, where the tree is given in Figure 10(a) and with $N=3200$ nodes and $p_{k}=144\,\beta^{5-k}\,\frac{\log{N}}{N}$. The results are averaged over 100 realizations, and error bars show the standard error but are typically smaller than the symbols.
C.1.3 Increased Number of Bottom Communities
We evaluate the accuracy of HCD algorithms on a BTSBM of depth 6, resulting in $K=2^{6}=64$ bottom communities. Figures 11 compare the performance of different HCD algorithms on BTSBM. We observe that bottom-up and top-down approaches perform similarly well, though one may outperform the other depending on $\beta$ and the chosen evaluation metric.
(a) AMI
(b) Tree similarity
(c) Estimated $K$
Figure 11: Performance of top-down and bottom-up algorithms on BTSBMs of depth 6, with $N=6400$ nodes ($K=64$ and 100 nodes per community) and $p_{k}=81\,\beta^{6-k}\,\frac{\log{N}}{N}$. The results are averaged over 10 realizations, and error bars show the standard error but are typically smaller than the symbols.
C.1.4 Another Bottom-up Approach
In this subsection, we consider another type of hierarchical community detection algorithm that first identifies bottom communities then aggregates them. The method called synthesis [fang2023t], differs from linkage-based approaches by employing the (sparse) neighbor-joining algorithm for aggregation. As a result, synthesis can handle more general tree structures than the binary rooted trees produced by bottom-up and top-down methods, and it outputs unrooted, potentially non-binary trees.
Figure 12 reports the performance of our focal bottom-up algorithm with synthesis. We evaluate performance using three metrics: bottom community accuracy, Robinson-Foulds distance from the ground-truth tree, and the recovery rate of the true hierarchical structures. The Robinson-Foulds distance is a metric to measure the distance between two potentially unrooted trees and is normalized to range from 0 to 1. The tree recovery rate is defined as the proportion of outputs with a Robinson-Foulds distance of zero. Although Fang and Rohe [fang2023t] use sparse neighbor-joining to accommodate non-binary trees, we apply the standard neighbor-joining algorithm for a fair comparison since our ground-truth tree structures are binary.
(a) Bottom accuracy
(b) Robinson-Foulds distance
(c) Recovery rate
Figure 12: Performance of synthesis and bottom-up algorithms on BTSBMs of depth 5, with $N=6400$ nodes ($K=32$ and 200 nodes per community) and $p_{k-1}/p_{k}=0.36$. The results are averaged over 10 realizations, and error bars show the standard error but are typically smaller than the symbols.
@@ C.2.1 Military Inter-alliance
We next consider the network of military alliances between countries. The data is provided by the Alliance Treaty Obligations and Provisions (ATOP) project [leeds2002alliance]. We select the year 2018 (as this is the most recent year available). We define two countries as allied if they share a defensive alliance (we do not consider non-aggression pacts, as those are more numerous and historically not necessarily well respected). This leads to a network of 133 countries and 1391 alliances. Some important countries such as India or Switzerland are missing as they do not share any defensive alliances with anybody. Moreover, the graph is not connected, as a small component made of three countries (China, North Korea and Cuba) is disconnected from the rest of the world.
Figures 13 and 14 show the output of bottom-up and top-down algorithms. Bottom-up predicts 7 bottom communities, which represent geopolitical alliances based on political affiliation and geography (European countries, Eurasian countries, Arabic countries, Western African countries and Central/Southern African countries). The top level splits the graph’s largest connected component into 3 clusters: Western countries, Eurasian countries, and African and Middle-East countries. While some of these clusters are also recovered by top-down HCD, the separation of African countries by top-down algorithm appears worse.
@@ C.2.2 Football Data Set
We also test HCD algorithms on the United States college (American) football dataset [girvan2002community]. This network represents the schedule of Division I games for the 2000 season. Each node in the network corresponds to a college football team, and edges represent the regular-season games between the teams. The teams are categorized into 11 conferences, in which games are more frequent between the members. Each conference has 8 to 12 teams. We exclude the ”independent” teams which do not belong to any conferences. Since the original community labels appear to be based on the 2001 season, while the edges represent the games played during the 2000 season, we proceed to the same correction as in [evans2010clique].